# Non-Convex Self-Concordant Functions: Practical Algorithms and Complexity Analysis

### v1

@@ 3.1 Regularized Newton’s Method (RNM)
The scheme is as follows,
$x_{j+1}=x_{j}-\tfrac{1}{1+\kappa\nu_{f,F}(x_{j})}(\nabla^{2}(f+F)(x_{j}))^{-1}\nabla f(x_{j}),\quad\textnormal{for all }j=0,1,2,\ldots$ (8)
[…]
Under Assumption 3.1, we derive the convergence guarantees in terms of $\nu_{f,F}(x)$ rather than gradient norm, which is insufficient to measure the stationarity due to the lack of Lipschitz gradients. Our results can be interpreted as a generalization of [Nesterov-2018-Lectures, Section 5.2] and $\nu_{f,F}(x)=\lambda_{f}(x)$ when $F=0$ (see [Nesterov-2018-Lectures, Page 354]). For functions that are weakly self-concordant, this measure relates to gradient norms of Moreau envelopes (see Appendix A.1). We summarize our results in the following theorems and defer the proofs to Section 4.
@@
If the Polyak-Łojasiewicz condition [Lojasiewicz-1963-Propriete, Polyak-1964-Gradient] holds at a limit point of $\{x_{j}\}_{j\geq 0}$, we show that the function value locally converges linearly in the following theorem.
@@
Suppose that Assumption 3.1 holds, $f+F$ is $\alpha$-strongly convex and the Polyak-Łojasiewicz condition holds true at an interior optimal solution $x^{\star}\in\mathrm{int}(\textbf{dom}(f))$ (i.e., there exists $\mu>0$ and a neighborhood $U$ of $x^{\star}$ in $\textbf{dom}(f)$ such that $\frac{1}{2}\|\nabla f(x)\|^{2}\geq\mu(f(x)-f(x^{\star}))$ for all $x\in U$). If $x^{\star}$ is a limit point of $\{x_{j}\}_{j\geq 0}$, we have $x_{j}\rightarrow x^{\star}$ and $f(x_{j})-f(x^{\star})$ converges $Q$-linearly to 0.
Remark 3.4.
The RNM scheme performs well for training convolutional neural networks (CNNs) but is less effective for nonnegative matrix factorization (NMF). An explanation lies in different choices and magnitudes of $F$. In our CNN experiments, $F$ is quadratic with the coefficients on the order of $10^{-2}$, which is consistent with the asymmetric Hessian spectra reported in neural network training [Zhang-2024-Transformers], where negative curvature is mild. In these settings, a small regularizer is sufficient to stabilize the updates and capture local curvature. In contrast, the objective function $f$ and its Hessian entries are much larger in scale for NMF. Near a local minimizer, applying the same globally chosen $F$ introduces excessive regularization, which distorts the update direction and ultimately degrades performance.
Algorithm 1 Adaptive Regularization Method
1:  Input: $0<\sigma_{\min}\leq\sigma_{0}$, $0<\eta_{1}\leq\eta_{2}<1$ and $0<\gamma_{1}<1<\gamma_{2}<\gamma_{3}$.
@@
4:   Compute $d_{j}\in\mathbb{R}^{n}$ and construct the model $m_{j}:\mathbb{R}_{+}\to\mathbb{R}\cup\{+\infty\}$ according to one of two options: (i) $d_{j}$ satisfies Eq. (9) and $m_{j}$ is defined in Eq. (10); and (ii) $d_{j}$ is defined in Eq. (17) and $m_{j}$ is defined in Eq. (18).
5:   if $\nabla f(x_{j})^{\top}d_{j}>0$ then
6:    $t_{j}\leftarrow 0$.
7:   else
8:    $t_{j}\leftarrow\mathop{\rm argmin}_{t\geq 0}m_{j}(t)$.
9:   end if
10:   Compute $r_{j}\leftarrow\frac{f(x_{j})-f(x_{j}+t_{j}d_{j})}{f(x_{j})-m_{j}(t_{j})}$.
@@
$x_{j+1}\leftarrow\begin{cases}x_{j}+t_{j}d_{j},&\textnormal{if }r_{j}\geq\eta_{1},\\ x_{j},&\textnormal{otherwise}.\end{cases}\quad\sigma_{j+1}\in\begin{cases}[\max(\sigma_{\min},\gamma_{1}\sigma_{j}),\sigma_{j}],&\textnormal{if }r_{j}\geq\eta_{2},\\ [\sigma_{j},\gamma_{2}\sigma_{j}],&\textnormal{if }r_{j}\in(\eta_{1},\eta_{2}),\\ [\gamma_{2}\sigma_{j},\gamma_{3}\sigma_{j}],&\mathrm{if}\ r_{j}\leq\eta_{1}.\end{cases}$
11:   if the termination condition depending on $(x_{j},d_{j},\sigma_{j})$ is satisfied then
12:    return $x_{j}$.
13:   end if
14:  end for
3.2 Adaptive Regularization Method (ARM)
@@
The scheme shares a similar spirit with the methods from [Cartis-2011-Adaptive, Cartis-2011-Adaptive2, Cartis-2022-Evaluation] and has been summarized in Algorithm 1. At each iteration, we compute a direction $d_{j}\in\mathbb{R}^{n}$ and a step-size $t_{j}$ by minimizing the model $m_{j}(t)$. Then, we compute the ratio $r_{j}$ to determine if a sufficient decrease is made. If $r_{j}\geq\eta_{1}$, we accept the update; otherwise, the iterate remains the same. Finally, we adjust the parameter $\sigma_{j}$ based on $r_{j}$ as described in Algorithm 1.
The key ingredients of Algorithm 1 consist of a specific rule for selecting the direction $d_{j}$. The first option can yield approximate first-order stationary points (see Theorem 3.5), while the second option, equipped with an oracle that detects negative curvature, can yield approximate second-order stationary points (see Theorem 3.8). In what follows, we provide all of details and convergence results.
First option: general descent directions
To begin, we derive a convergence result for general update directions $d_{j}$. The only requirement is that $d_{j}=d(x_{0},\ldots,x_{j},\sigma_{j})$ is a descent direction when the parameter $\sigma_{j}$ is sufficiently large:
$\exists\bar{\sigma}>0,\textnormal{ s.t. }\nabla f(x)^{\top}d_{j}\leq 0\textnormal{ if }\sigma_{j}\geq\bar{\sigma}.$ (9)
Under the same notation of Proposition 2.6, we view $f$ as $\sigma_{j}F$-based $\kappa$-self-concordant and define the local model $m_{j}(t)$ as follows,
@@
$m_{j}(t)=\begin{cases}f(x_{j})-\rho_{j}t+\kappa^{-2}\,\omega_{\star}(\kappa t\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}),&\textnormal{if }\delta_{j}+\sigma_{j}\Delta_{j}\geq 0,\\ 0,&\mathrm{if}\ \delta_{j}+\sigma_{j}\Delta_{j}<0\textnormal{ and }t=0,\\ +\infty,&\textnormal{otherwise}.\end{cases}$ (10)
Thus, we only compute $t_{j}$ when $\delta_{j}+\sigma_{j}\Delta_{j}\geq 0$ and $\rho_{j}\geq 0$. In this case, minimizing Eq. (10) with respect to $t$ yields the solution in the form of
$t_{j}=\tfrac{\rho_{j}}{\delta_{j}+\sigma_{j}\Delta_{j}+\kappa\rho_{j}\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}}.$ (11)
@@
Suppose that Assumption 3.1 holds and Eq. (9) is satisfied. Let $m_{j}$ be defined in Eq. (10), and the termination condition be
$\tfrac{\nabla f(x_{j})^{\top}d_{j}}{\sqrt{\nabla^{2}(f+\sigma_{j}F)(x_{j})[d_{j},d_{j}]}}\leq\epsilon.$ (12)
Then, we have $\sigma_{j}\leq\max(\gamma_{3}\bar{\sigma},\sigma_{0},\gamma_{3})$ for all $j$, and Algorithm 1 terminates within at most $O(\epsilon^{-2})$ iterations.
Note that it is possible to terminate at a point $x_{j}$ where both $\nabla f(x_{j})$ and $\nabla^{2}(f+\sigma_{j}F)$ have large magnitudes. To mitigate this issue, we introduce two specific choices for $d_{j}$ and refine the termination conditions accordingly.
The first choice is preconditioned gradient descent. Indeed, we define
$d_{j}=-H_{j}\nabla f(x),\quad H_{j}\succ 0,$ (13)
where $H_{j}$ is the preconditioner that depends on $x_{0},\ldots,x_{j}$. This framework covers both gradient descent ($H_{j}=I$) and quasi-Newton methods. Since $H_{j}\succ 0$, we have $\nabla f(x_{j})^{\top}d_{j}\leq 0$, ensuring that Eq. (9) is satisfied.
Corollary 3.6.
@@
Suppose that Assumption 3.1 holds, and there exists $\Lambda>0$ such that $\|H_{j}^{1/2}\nabla^{2}(f+\sigma_{j}F)H_{j}^{1/2}\|_{\mathrm{op}}\leq\Lambda$ for all $j$. Let $d_{j}$ and $m_{j}$ be defined in Eq. (13) and Eq. (10) and the termination condition be $\sqrt{\nabla f(x_{j})^{\top}H_{j}\nabla f(x_{j})}\leq\epsilon$. Then, we have $\sigma_{j}\leq\max(\sigma_{0},\gamma_{3})$, and Algorithm 1 terminates within at most $O(\epsilon^{-2})$ iterations.
The second choice is adaptive regularized Newton’s method. Indeed, we define
$d_{j}=-(\nabla^{2}(f+\sigma_{j}F)(x_{j}))^{-1}\nabla f(x_{j}),$ (14)
@@
The termination condition in Eq. (12) becomes $\nu_{f,\sigma_{j}F}(x_{j})\leq\epsilon$. If $\delta_{j}+\sigma_{j}\Delta_{j}\geq 0$ and $\rho_{j}\geq 0$, the optimal step size is
@@
Corollary 3.7.
Suppose that Assumption 3.1 holds. Let $d_{j}$ and $m_{j}$ be defined in Eq. (14) and Eq. (15), and the termination condition be $\nu_{f,\sigma_{j}F}(x_{j})\leq\epsilon$. Then, we have Algorithm 1 terminates within at most $O(\epsilon^{-2})$ iterations.
Second option: negative curvature descent

### v2

@@ 3.1 Regularized Newton’s Method (RNM)
At each iteration, Algorithm 1 (RNM) computes a direction $d_{j}\in\mathbb{R}^{n}$ that minimizes the right-hand side of inequality (6). This is equivalent to maximizing $\eta_{j}=\tfrac{-\nabla f(x_{j})^{\top}d_{j}}{\sqrt{d_{j}^{\top}(\nabla^{2}(f+F)(x_{j}))d_{j}}}$ w.r.t. $d_{j}$ and yields Algorithm 1.
Algorithm 1 Regularized Newton’s Method
1:Initialization: $x_{0}\in\textbf{dom}(f)$.
2:for $j=0,1,2,\ldots$ do
3:  $d_{j}\leftarrow-(\nabla^{2}(f+F)(x_{j}))^{-1}\nabla f(x_{j})$.
4:  $\lambda_{j}\leftarrow\sqrt{-\nabla f(x_{j})^{\top}d_{j}}$.
5:  $t_{j}\leftarrow\tfrac{1}{1+\kappa\lambda_{j}}$.
6:  $x_{j+1}\leftarrow x_{j}+t_{j}d_{j}$.
7:end for
Let $A^{+}$ denote the pseudoinverse of $A\in\mathbb{R}^{n\times n}$ and $\sqrt{x}=\infty$ if $x<0$. These conventions will be convenient in Algorithm 2. To analyze Algorithm 1, we define the square root of regularized Newton decrement as
$\nu_{f,F}(x_{j})=\sqrt{\delta_{j}+\Delta_{j}},\;\;\mathrm{where}\ d_{j}=-(\nabla^{2}(f+F)(x_{j}))^{+}\nabla f(x_{j}),$ (8)
Under Assumption 3.1, we derive convergence guarantees in terms of $\nu_{f,F}(x)$ rather than the gradient norm, which is insufficient to measure the stationarity due to the lack of Lipschitz gradients. Our results can be interpreted as a generalization of [58, Section 5.2] and $\nu_{f,F}(x)=\lambda_{f}(x)$ when $F=0$ (see [58, Page 354]). For functions that are weakly self-concordant, this measure relates to gradient norms of Moreau envelopes (see Appendix A.1). We summarize our results in the following two theorems.
@@
Proof.
Proof Proposition 2.6 guarantees that $x_{j}\in\textbf{dom}(f)$, $f(x_{j})\leq f(x_{0})$ for all $j\geq 0$ and $f(x_{j})-f(x_{j+1})\geq\frac{(\nu_{f,F}(x_{j}))^{2}}{2(1+\Gamma_{f}(x_{0}))}$. Summing this inequality over $j=0,1,2,\ldots$ yields
$f(x_{0})-\inf_{x\in\mathbb{R}^{n}}f(x)\geq\sum_{j=0}^{k}f(x_{j})-f(x_{j+1})\geq\sum_{j=0}^{k}\tfrac{\nu_{f,F}(x_{j})^{2}}{2(1+\Gamma_{f}(x_{0}))}\geq\tfrac{(k+1)(\min_{0\leq j\leq k}\nu_{f,F}(x_{j}))^{2}}{2(1+\Gamma_{f}(x_{0}))}.$
This completes the proof. $\square$
If the Polyak-Łojasiewicz (PL) condition [42, 62] holds in a neighborhood of a limit point of $\{x_{j}\}_{j\geq 0}$, we further have
@@
Suppose that Assumption 3.1 holds, $f+F$ is $\alpha$-strongly convex and $f$ satisfies the Polyak-Łojasiewicz (PL) condition with constant $\mu>0$ in a neighborhood $U$ in $\textbf{dom}(f)$ of an optimal solution $x^{\star}$, (i.e., $\frac{1}{2}\|\nabla f(x)\|^{2}\geq\mu(f(x)-f(x^{\star}))$ for all $x\in U$). If $x^{\star}$ is a limit point of $\{x_{j}\}_{j\geq 0}$, generated by Algorithm 1, we have $x_{j}\rightarrow x^{\star}$ and $f(x_{j})-f(x^{\star})$ converges $Q$-linearly to 0.
Proof.
Proof In [2, Proposition 3.3 and Theorem 3.4], it is proved that if $f$ satisfies the PL condition in a neighborhood $U$ of $x_{\star}$ then $x_{j}\to x_{\star}$ if there exist $c_{1},c_{2}>0$ such that $f(x_{j})-f(x_{j+1})\geq c_{1}\|\nabla f(x_{j})\|^{2}$ and $\|\nabla f(x_{j})\|\geq c_{2}\|x_{j}-x_{j+1}\|$ for all $x_{j},x_{j+1}\in U$. Now by Proposition 2.6 and the facts that $f(x_{j})\leq f(x_{0})$ and $\|\nabla^{2}(f+F)(x_{j})\|_{\mathrm{op}}\leq L$, for some $L>0$, for all $x_{j}\in U$, we have
We also have $\|\nabla f(x_{j})\|\geq\alpha\|x_{j+1}-x_{j}\|$. Hence, from the results in [2] stated above, it follows that, $x_{j}\to x_{\star}$.
It remains to prove the linear convergence of function values. Since $x_{j}\to x^{\star}$, there exists $K\in\mathbb{N}$ such that $x_{j}\in U$ for all $j\geq K$. Applying Eq. (9) and the PL inequality yields $f(x_{j})-f(x_{j+1})\geq\frac{\mu}{(1+\Gamma_{f}(x_{0}))L}(f(x_{j})-f(x^{\star}))$ for all $j\geq K$. This implies that
$f(x_{j+1})-f(x^{\star})\leq\left(1-\tfrac{\mu}{(1+\Gamma_{f}(x_{0}))L}\right)(f(x_{j})-f(x^{\star})),$
which implies that $f(x_{j})-f(x^{\star})$ converges $Q$-linearly to 0. $\square$
3.2 Adaptive Regularization Method (ARM)
Algorithm 2 Adaptive Regularization Method (ARM)
1:Input: $0<\sigma_{\min}\leq\sigma_{0}$, $0<\eta_{1}\leq\eta_{2}<1$, $0<\gamma_{1}<1<\gamma_{2}<\gamma_{3}$. Opt = 1 or 2.
@@ 3.2 Adaptive Regularization Method (ARM)
4:  Opt=1: Given $H_{j}$, compute $d_{j}$ and $m_{j}$ by Eq. (10), $\rho_{j}\leftarrow-\nabla f(x_{j})^{\top}d_{j}$, $\delta_{j}\leftarrow d_{j}^{\top}\nabla^{2}f(x_{j})d_{j}$, and $\Delta_{j}\leftarrow d_{j}^{\top}\nabla^{2}F(x_{j})d_{j}$. return $x_{j}$ if ${\rho_{j}}\leq\epsilon\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}$ and $\rho_{j}\geq 0$ and $\delta_{j}+\sigma_{j}\Delta_{j}\geq 0$.
5:  Opt=2: Compute $v_{j}$ s.t. $\|v_{j}\|=1$ and $\nabla^{2}f(x_{j})v_{j}=\lambda_{\min}(\nabla^{2}f(x_{j}))v_{j}$, $d_{j}$ by (14), $m_{j}$ by (15), $\rho_{j}\leftarrow-\nabla f(x_{j})^{\top}d_{j}$, $\delta_{j}\leftarrow d_{j}^{\top}\nabla^{2}f(x_{j})d_{j}$, and $\Delta_{j}\leftarrow d_{j}^{\top}\nabla^{2}F(x_{j})d_{j}$. return $x_{j}$ if $\nu_{f,\sigma_{j}F}(x_{j})\leq\epsilon_{g}$ and $\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\sigma_{j}\sqrt{\epsilon_{H}}v_{j}^{\top}\nabla^{2}F(x_{j})v_{j}$.
6:  $t_{j}\leftarrow\mathop{\rm argmin}_{t\geq 0}m_{j}(t)$, $t_{j}\leftarrow 0$ if $t_{j}=\infty$ or any tie occurs.
7:  $r_{j}\leftarrow\frac{f(x_{j})-f(x_{j}+t_{j}d_{j})}{f(x_{j})-m_{j}(t_{j})}$, where $f(x_{j}+t_{j}d_{j})\leftarrow+\infty$ if $x_{j}+t_{j}d_{j}\notin\textbf{dom}(f)$ and $\frac{0}{0}=0$.
@@
$x_{j+1}\leftarrow\begin{cases}x_{j}+t_{j}d_{j},&\textnormal{if }r_{j}>\eta_{1},\\ x_{j},&\textnormal{otherwise}.\end{cases}\quad\sigma_{j+1}\in\begin{cases}[\max(\sigma_{\min},\gamma_{1}\sigma_{j}),\sigma_{j}],&\textnormal{if }r_{j}\geq\eta_{2},\\ [\sigma_{j},\gamma_{2}\sigma_{j}],&\textnormal{if }r_{j}\in(\eta_{1},\eta_{2}),\\ [\gamma_{2}\sigma_{j},\gamma_{3}\sigma_{j}],&\mathrm{if}\ r_{j}\leq\eta_{1}.\end{cases}$
8:end for
@@
Our second algorithm Algorithm 2 (ARM) is similar in spirit to the methods in [11, 12, 13]. At each iteration, Algorithm 2 computes a direction $d_{j}\in\mathbb{R}^{n}$ and a step-size $t_{j}$ by minimizing the model $m_{j}(t)$. It then computes the ratio $r_{j}$ to determine if a sufficient decrease is made. If $r_{j}>\eta_{1}$, it accepts the update; otherwise, the iterate is not updated. Finally, it adjusts the parameter $\sigma_{j}$ based on $r_{j}$.
Algorithm 2 has two options (Opt =1=1 or $2$) for selecting the search direction $d_{j}$ at every iteration. The first option is preconditioned gradient descent, given by
@@
$d_{j}=-H_{j}\nabla f(x_{j}),\quad m_{j}:\mathbb{R}_{+}\to\mathbb{R},m_{j}(t)=f(x_{j})-\rho_{j}t+\kappa^{-2}\,\omega_{\star}(\kappa t\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}),$ (10)
The convergence of option 1 of ARM depends on the choice of $H_{j}$. Our convergence analysis (Theorem 3.4 and Lemma 3.5) covers gradient descent ($H_{j}=I$), quasi-Newton methods, and Newton-type methods. Under the assumptions of Lemmas 3.5 or 3.6, we are able to show that option 1 of ARM returns an approximate first-order stationary point. The second option of ARM, which requires an oracle that detects negative curvature, returns an approximate second-order stationary point (see Theorem 3.7). In what follows, we provide additional details and convergence results.
@@
Suppose that Assumption 3.1 holds. If there exists $\sigma_{\max}>0$ such that $\sigma_{j}\leq\sigma_{\max}$ for all $j\geq 0$, then Algorithm 2 (ARM) with Opt =1=1 terminates within at most $O(\epsilon^{-2})$ iterations.
Proof.
Proof Fix an integer $k\geq 1$ and suppose Algorithm 2 has not terminated before iteration $k$. Let $\mathcal{S}_{k}$ and $\mathcal{U}_{k}$ denote the sets of successful and unsuccessful iterations up to $k$ in Algorithm 2,
$\mathcal{S}_{k}:=\{0\leq j\leq k-1:\ r_{j}>\eta_{1}\},\qquad\mathcal{U}_{k}:=\{0\leq j\leq k-1:\ r_{j}\leq\eta_{1}\}.$
For each $j\in\mathcal{S}_{k}$, we have $x_{j+1}=x_{j}+t_{j}d_{j}$ and therefore
$f(x_{j})-f(x_{j+1})=r_{j}\bigl(f(x_{j})-m_{j}(t_{j})\bigr)\geq\eta_{1}\bigl(f(x_{j})-m_{j}(t_{j})\bigr).$
Hence, if there exists a constant $\widetilde{\epsilon}>0$ such that $f(x_{j})-m_{j}(t_{j})\geq\widetilde{\epsilon}$ for all $j\in\mathcal{S}_{k}$, then
$f(x_{0})-\inf_{z\in\mathbb{R}^{n}}f(z)\geq\sum_{j\in\mathcal{S}_{k}}\bigl(f(x_{j})-f(x_{j+1})\bigr)\geq\eta_{1}\,\lvert\mathcal{S}_{k}\rvert\,\widetilde{\epsilon}$,
which implies $\lvert\mathcal{S}_{k}\rvert=O(\widetilde{\epsilon}^{-1})$.
We now lower bound the model decrease on successful iterations. For option 1 in Algorithm 2, whenever $j\in\mathcal{S}_{k}$, a direct calculation from the explicit form of $m_{j}$ yields the bound
$f(x_{j})-m_{j}(t_{j})\ \geq\ \kappa^{-2}\omega\left(\tfrac{-\kappa\nabla f(x_{j})^{\top}d_{j}}{\sqrt{\nabla^{2}(f+\sigma_{j}F)(x_{j})[d_{j},d_{j}]}}\right).$
Since the algorithm has not terminated, the option 1 termination condition is violated at $x_{j}$, i.e.,
$\tfrac{\rho_{j}}{\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}}=\tfrac{-\nabla f(x_{j})^{\top}d_{j}}{\sqrt{\nabla^{2}(f+\sigma_{j}F)(x_{j})[d_{j},d_{j}]}}>\epsilon.$
Consequently, for all $j\in\mathcal{S}_{k}$,
$f(x_{j})-m_{j}(t_{j})>\kappa^{-2}\omega(\kappa\epsilon),$
and hence $\lvert\mathcal{S}_{k}\rvert=O(\epsilon^{-2})$.
Finally, we bound the number of unsuccessful iterations using the update rule for $\sigma_{j}$. For any $j\in\mathcal{S}_{k}$ we have $\sigma_{j+1}\geq\max(\sigma_{\min},\gamma_{1}\sigma_{j})\geq\gamma_{1}\sigma_{j}$, while for any $j\in\mathcal{U}_{k}$ we have $\sigma_{j+1}\geq\gamma_{2}\sigma_{j}$. Thus,
$\sigma_{0}\,\gamma_{1}^{\lvert\mathcal{S}_{k}\rvert}\gamma_{2}^{\lvert\mathcal{U}_{k}\rvert}\ \leq\ \sigma_{k}\ \leq\ \sigma_{\max}.$
Taking logarithms and rearranging yields
$\lvert\mathcal{U}_{k}\rvert\leq\tfrac{\log(\sigma_{\max}/\sigma_{0})-\lvert\mathcal{S}_{k}\rvert\log(\gamma_{1})}{\log(\gamma_{2})},$
and therefore $k=\lvert\mathcal{S}_{k}\rvert+\lvert\mathcal{U}_{k}\rvert=O(\lvert\mathcal{S}_{k}\rvert)=O(\epsilon^{-2})$. $\square$
In many practical settings the preconditioner satisfies $H_{j}\succ 0$, e.g., gradient descent with $H_{j}=I$, quasi-Newton methods [41], and inverse Gauss–Newton approximations. Under this additional structure, a spectral bound on the preconditioned Hessian ensures that $\sigma_{j}$ remains uniformly bounded, as shown in Lemma 3.5 below, so the $O(\epsilon^{-2})$ complexity of Theorem 3.4 applies directly.
Lemma 3.5.
Suppose that Assumption 3.1 holds and $H_{j}\succ 0$. Then, we have $\sigma_{j}\leq\max(\sigma_{0},\gamma_{3})$ for ARM with Opt =1=1.
Proof.
Proof With $d_{j}=-H_{j}\nabla f(x_{j})$ and $H_{j}\succ 0$, we have $\nabla f(x_{j})^{\top}d_{j}=-\nabla f(x_{j})^{\top}H_{j}\nabla f(x_{j})\leq 0$ for all $j$. Suppose for contradiction that there exists an index $j$ such that $\sigma_{j+1}>\max(\sigma_{0},\gamma_{3})\geq\sigma_{j}$. Then $\sigma_{j}\geq\sigma_{j+1}/\gamma_{3}\geq 1$. Assumption 3.1 implies $\delta_{j}+\Delta_{j}\geq 0$, and hence $\delta_{j}+\sigma_{j}\Delta_{j}\geq 0$. By Proposition 2.6, we have $x_{j}+t_{j}d_{j}\in\textbf{dom}(f)$ and
$f(x_{j})-f(x_{j}+t_{j}d_{j})\ \geq\ \rho_{j}t_{j}-\kappa^{-2}\omega_{\star}\!\bigl(\kappa t_{j}\sqrt{\delta_{j}+\Delta_{j}}\bigr).$
On the other hand, by the definition of the option 1 model $m_{j}$,
$f(x_{j})-m_{j}(t_{j})=\rho_{j}t_{j}-\kappa^{-2}\omega_{\star}\!\bigl(\kappa t_{j}\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}\bigr).$
Since $\sigma_{j}\geq 1$ and $\omega_{\star}$ is increasing on $\mathbb{R}_{+}$, we obtain $f(x_{j})-f(x_{j}+t_{j}d_{j})\geq f(x_{j})-m_{j}(t_{j})$, so $r_{j}\geq 1>\eta_{2}$. Algorithm 2 would therefore enforce $\sigma_{j+1}\leq\sigma_{j}$, contradicting $\sigma_{j+1}>\sigma_{j}$. Hence, $\sigma_{j}\leq\max(\sigma_{0},\gamma_{3})$ for all $j$. $\square$
We can further specialize ARM to obtain an adaptive method by leveraging the self-concordant structure of $f+\sigma_{j}F$ for $\sigma_{j}\geq 1$. In particular, we define
$d_{j}=-(\nabla^{2}(f+\sigma_{j}F)(x_{j}))^{+}\nabla f(x_{j}),$ (11)
@@
The termination condition becomes $\nu_{f,\sigma_{j}F}(x_{j})\leq\epsilon$. If $\delta_{j}+\sigma_{j}\Delta_{j}>0$ and $\rho_{j}\geq 0$, the optimal step size is
@@
Lemma 3.6.
Suppose that Assumption 3.1 holds. Let $d_{j}$ be defined by Eq. (11). Then, we have $\sigma_{j}\leq\max(\sigma_{0},1,\gamma_{3})$ for ARM with Opt =1=1.
Proof.
Proof Suppose for contradiction that there exists $j$ such that $\sigma_{j+1}>\max(\sigma_{0},1,\gamma_{3})\geq\sigma_{j}$. Since $\sigma_{j+1}>\gamma_{3}$ and Algorithm 2 always enforces $\sigma_{j+1}\leq\gamma_{3}\sigma_{j}$, we have $\sigma_{j}\geq\frac{\sigma_{j+1}}{\gamma_{3}}>1$. Let $\rho_{j}:=-\nabla f(x_{j})^{\top}d_{j}\geq 0$, $\delta_{j}:=\nabla^{2}f(x_{j})[d_{j},d_{j}]$, and $\Delta_{j}:=\nabla^{2}F(x_{j})[d_{j},d_{j}]\geq 0$. Assumption 3.1 implies $\delta_{j}+\Delta_{j}\geq 0$ and hence $\delta_{j}+\sigma_{j}\Delta_{j}\geq 0$. By Proposition 2.6, we have $x_{j}+t_{j}d_{j}\in\textbf{dom}(f)$ and
$f(x_{j})-f(x_{j}+t_{j}d_{j})\ \geq\ \rho_{j}t_{j}-\kappa^{-2}\omega_{\star}\!\bigl(\kappa t_{j}\sqrt{\delta_{j}+\Delta_{j}}\bigr).$
On the other hand, since $d_{j}=-(\nabla^{2}(f+\sigma_{j}F)(x_{j}))^{+}\nabla f(x_{j})$, we have from (8) that $\rho_{j}=\nabla f(x_{j})^{\top}\bigl(\nabla^{2}(f+\sigma_{j}F)(x_{j})\bigr)^{+}\nabla f(x_{j})=\bigl(\nu_{f,\sigma_{j}F}(x_{j})\bigr)^{2}=\delta_{j}+\sigma_{j}\Delta_{j}$, and hence from (12) that
$f(x_{j})-m_{j}(t_{j})=\rho_{j}t_{j}-\kappa^{-2}\omega_{\star}\!\bigl(\kappa t_{j}\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}\bigr).$
Because $\sigma_{j}\geq 1$ and $\omega_{\star}$ is increasing on its domain, we have $\omega_{\star}\!\bigl(\kappa t_{j}\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}\bigr)\geq\omega_{\star}\!\bigl(\kappa t_{j}\sqrt{\delta_{j}+\Delta_{j}}\bigr)$; therefore $f(x_{j})-f(x_{j}+t_{j}d_{j})\geq f(x_{j})-m_{j}(t_{j})$, i.e., $r_{j}\geq 1>\eta_{2}$. By the update rule in Algorithm 2, $r_{j}\geq\eta_{2}$ forces $\sigma_{j+1}\leq\sigma_{j}$, contradicting $\sigma_{j+1}>\sigma_{j}$. $\square$